# Unlocking Adversarial Suffix Optimization Without Affirmative Phrases: Efficient Black-box Jailbreaking via LLM as Optimizer

### v1

@@ Abstract
Despite prior safety alignment efforts, mainstream LLMs can still generate harmful and unethical content when subjected to jailbreaking attacks. Existing jailbreaking methods fall into two main categories: template-based and optimization-based methods. The former requires significant manual effort and domain knowledge, while the latter, exemplified by Greedy Coordinate Gradient (GCG), which seeks to maximize the likelihood of harmful LLM outputs through token-level optimization, also encounters several limitations: requiring white-box access, necessitating pre-constructed affirmative phrase, and suffering from low efficiency. In this paper, we present ECLIPSE, a novel and efficient black-box jailbreaking method utilizing optimizable suffixes. Drawing inspiration from LLMs’ powerful generation and optimization capabilities, we employ task prompts to translate jailbreaking goals into natural language instructions. This guides the LLM to generate adversarial suffixes for malicious queries. In particular, a harmfulness scorer provides continuous feedback, enabling LLM self-reflection and iterative optimization to autonomously and efficiently produce effective suffixes. Experimental results demonstrate that ECLIPSE achieves an average attack success rate (ASR) of 0.92 across three open-source LLMs and GPT-3.5-Turbo, significantly surpassing GCG in 2.4 times. Moreover, ECLIPSE is on par with template-based methods in ASR while offering superior attack efficiency, reducing the average attack overhead by 83%.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have made significant strides with represented by ChatGPT [1], Claude [2], and Gemini [3], demonstrating remarkable capabilities in intelligent question answering [4], code generation [5], task planning [6], and logical reasoning [7]. These advancements are bringing profound changes to the way people live and work. Mainstream LLM products undergo safety alignment processes before deployment to ensure their outputs align with human moral values and laws. However, recent research [8, 9, 10, 11] reveals that despite these preventive measures, LLMs remain vulnerable to jailbreaking attacks. Attackers are capable of crafting sophisticated prompts that manipulate LLMs into harmful responses, posing significant threats to their safety and credibility.
Existing jailbreaking methods primarily fall into two categories: template-based and optimization-based methods. The template-based methods leverage patterns derived from successful jailbreak hints [8, 9] or incorporate insights from psychology and social engineering [10, 12] to devise effective jailbreak templates, combining manual and automated approaches. While insightful, they require significant manual effort and domain knowledge, limiting the number of candidate prompts. They also often depend on the target LLM’s ability to understand and follow specific instructions in their well-crafted jailbreak prompts. Conversely, optimization-based methods treat jailbreaking as a discrete optimization problem, searching for token combinations that maximize the likelihood of specific malicious responses [11, 13, 14]. Greedy Coordinate Gradient (GCG)  [11] is a notable example, optimizing suffixes to prompt predetermined affirmative phrases to induce the target LLM to continue outputting malicious content. Although these methods are adaptable and offer extensive candidate generation but require white-box LLM access and manually design affirmative phrases as optimization targets, often leading to suboptimal effectiveness and efficiency, as shown in Table 1.
@@
In this paper, we propose ECLIPSE, an efficient black-box jailbreaking method with optimizable suffixes, for exploiting the vulnerabilities of LLMs. Our key insight is that LLMs can be employed as both generators and optimizers in the jailbreaking process, iteratively refining their outputs based on feedback to ultimately achieve successful attacks, which is inspired by recent studies [15]. More specifically, we make an observation: it is feasible to articulate the goal of generating jailbreaking suffixes in natural language, prompting the LLM to produce candidate suffixes. By providing appropriate feedback on each suffix, we can guide the LLM to iteratively refine its outputs, culminating in successful jailbreaking suffixes. Building on this foundation, we develop ECLIPSE, which encapsulates this process. Firstly, we introduce a novel task-prompting strategy for LLMs, where they engage in role-playing to generate suffixes that perturb the hidden space features and persuade the counterpart chatbot to respond to a malicious query, thereby aligning the LLM with our specific jailbreaking objectives. Secondly, to evaluate the efficacy of these generated suffixes, we engage an automated harmfulness scorer, which provides quantitative assessments. Furthermore, by maintaining a historical record of generated suffixes and their efficacy scores, we provide essential references that enable the attacker LLM to self-reflect and optimize its solutions. Our experimental evaluations, conducted on three open-source models and GPT-3.5-Turbo, demonstrate that ECLIPSE achieves an average Attack Success Rate (ASR) of 0.92, surpassing existing optimization-based methods (i.e., GCG) in effectiveness, efficiency, and naturalness. Additionally, ECLIPSE exhibits a comparable ASR to template-based methods while significantly reducing the average attack time overhead by 83% and the number of queries by 45%, thereby enhancing its efficiency dramatically. Our implementation is available at https://github.com/lenijwp/ECLIPSE.
@@ 2.1 Related Work
Template-based Methods. Template-based methods leverage patterns derived from successful jailbreak hints [8, 9] or incorporate insights from psychology and social engineering [10, 12] to devise effective jailbreak templates, either manually or automatically. DeepInception [10] introduces manually crafted templates that embed nested scenarios to hypnotize LLMs into producing malicious outputs. RED-EVAL [16] employs a Chain of Utterances (CoU)-based prompt to extract harmful information step by step. GPTFUZZER [9], employs software fuzzing test principles, to collect a variety of successful jailbreaking templates as seeds and apply mutations to spawn novel prompts. Masterkey [8] fine-tunes an LLM using reinforcement learning on collected workable jailbreak prompts to unearth underlying jailbreak mechanisms and learn potent patterns autonomously. PAIR [12] enables LLMs to autonomously generate and refine jailbreaking prompts by providing some pre-designed general strategies such as obfuscating sensitive words and engaging in role-playing scenarios. Additionally, some studies convert jailbreaking prompts into encrypted or low-resource languages [17, 18], exploiting the limited alignment in these languages to facilitate jailbreaking. These methods can efficiently achieve jailbreaking by directly black-box queries to the target LLM’s API, even obviating the need for multiple rounds of iterative trials.
Optimization-based Methods. Optimization-based methods are commonly employed for generating adversarial examples in NLP tasks, particularly in discriminative tasks [19, 20, 21]. These methods typically model attack targets by manipulating embeddings or predicting logits, thereby facilitating the gradient-based optimization search for candidate tokens. Among these, the Greedy Coordinate Gradient (GCG) [11] stands out as a seminal work focused on the optimization-based jailbreaking of generative LLMs. GCG aims to craft adversarial suffixes that prompt LLMs to specific maliciously affirmative phrases (e.g., Sure, here is …). The key insight behind this is that if the LLM’s response begins with affirmative phrases, there is a high probability that it will continue to generate more malicious content. GCG streamlines the suffix generation process by combining greedy and gradient-based discrete optimization. In the GCG framework, a malicious query is represented by a sequence of $n$ tokens, $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$. The aim is to identify an optimizable suffix $\mathbf{s}=(s_{1},s_{2},\ldots,s_{m})$ that, when concatenated to $\mathbf{x}$, maximizes the probability of eliciting the predefined affirmative phrase sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{k})$ from the target LLM. The optimization goal is formally defined as:
@@
$P(\mathbf{y}\mid\mathbf{x}\oplus\mathbf{s})=\prod_{i=1}^{k}P(y_{i}\mid\mathbf{x}\oplus\mathbf{s},y_{1},y_{2},\ldots,y_{i-1})$ (2)
where $\oplus$ denotes the concatenation operation. Recently, some studies have aimed to further enhance token optimization methods. Notably, RIPPLE [13] proposes replacing affirmative phrases with subconscious exploitations, aiming to refine and streamline the optimization process. Similarly, AutoDAN [14] introduces a hierarchical genetic algorithm for optimizing discrete tokens, effectively bypassing traditional gradient propagation techniques.
@@ 2.2 Motivation
Template-based methods necessitate a significant amount of manual effort to design, collect, and even tailor templates for each specific malicious query (e.g., DeepInception [10]). The number of jailbreaking prompts that can be filled and combined based on a specific set of templates is also relatively limited. Moreover, the efficacy of many templates relies heavily on the instruction-following capabilities of LLMs [22], which constrains the universal applicability of the method. In contrast, optimization-based methods have a lower dependence on the instruction-following abilities of LLMs and are not easily neutralized by alignment training [23]. However, they also face several challenges. Optimization methods such as GCG typically necessitate presetting expected target responses, often as affirmative phrases. This requirement introduces two primary challenges. First, the discrete nature of LLM tokens makes optimizing input sequences for NLP models particularly difficult. Ad hoc selection of affirmative phrases can lead to optimization goals that are hard to achieve. Secondly, while GCG posits that if an LLM starts its response with affirmative phrases, it is likely to continue to respond positively, recent research [13] suggests this may not always be the case. LLMs might start with affirmative phrases but subsequently refuse to comply. Furthermore, the inefficiency of the optimization process and the need for white-box access significantly limit its practical usage in real-world scenarios. Hence, the goal of this paper is to explore an efficient black-box suffix optimization approach to achieve jailbreaking, without specific affirative phrases.
@@ 3.1 Threat Model
We adhere to the threat model initially proposed by GCG [11], which involves finding a specific suffix for a given input to induce the target LLM to output harmful or unethical responses. However, we extend this model to more accurately reflect real-world scenarios. Our primary objective is to transcend the limitations of predefined affirmative phrases while still facilitating the optimization of effective jailbreaking suffixes. Furthermore, we aim to transition from a white-box scenario, where attackers have access to the model’s weights and gradients, to a black-box scenario. This shift is motivated by the prevalent deployment of advanced LLMs as API services, where direct access to the model’s internals is not available, and attackers can only interact through API calls and receive textual responses.
@@ 3.2 Our Intituitive Idea
Recently, employing LLMs as black-box optimizers for complex objective functions represents a cutting-edge approach [24, 15, 25, 26]. These LLMs engage in dialogical interactions to progressively formulate recommendations, thereby incrementally refining their outputs guided by predefined optimization goals and continuous feedback. Notably effective in managing optimization tasks articulated through natural language, this method has extended the utility of LLMs beyond numeric problems to include complex tasks like coding and text generation [27].
The demonstrated capability of LLMs to adapt and refine complex, non-numeric tasks through dialogical interactions and feedback-driven optimization raises an intriguing question: Can the generation and optimization capabilities of LLMs be harnessed to automatically optimize adversarial suffixes for effective jailbreaking? Recent research, such as the PAIR [12], has shown that LLMs can indeed refine jailbreak prompts by devising rewriting and storytelling strategies. Our study, however, delves into a more specialized area: examining whether LLMs can explicitly generate and optimize suffixes to meet the specific goals of jailbreaking. This particular focus presents a rigorous challenge to their optimization capabilities, testing their ability to adapt and perform under narrowly defined constraints. In our LLaMA2 prototype depicted in  Figure 1, we use an LLM as an attacker to generate and refine suffixes aimed at manipulating a chatbot to respond to malicious queries. We iteratively test these suffixes, concatenate them to the original queries, and feed the results back to the LLM to foster the generation of new, more effective candidates. Through iterative testing and feedback, the attacker LLM demonstrates two critical capabilities in identifying successful jailbreaking suffixes. Our findings highlight two critical capabilities of LLMs in generating and optimizing jailbreaking suffixes. Firstly, LLMs demonstrate a promising optimization aptitude, requiring only quantifiable feedback to self-reflect and refine their outputs efficiently. Secondly, their strong generation capacity allows them to produce numerous candidate suffixes that are semantically clearer and more natural than those generated by methods like GCG. Meanwhile, LLM-generated suffixes show a high correlation with given malicious queries, indicating that the LLM comprehends and incorporates multiple potential responses to the posed questions. These knowledge-enriched suffixes are more likely to prompt the target LLM to reveal elements of its memorization for jailbreaking.
@@ 3.3 Method Design
Inspired by the insights from leveraging LLMs as powerful optimizers, we propose a novel and efficient black-box jailbreaking method with optimizable suffixes, ECLIPSE. Following established workflows for LLM optimizers [15, 24], ECLIPSE streamlines the suffix generation process for jailbreaking. The method starts with task prompting, translating our jailbreaking objectives into natural language for the attacker LLM. We employ an automated tool to score the harmfulness of generated suffixes, providing continuous feedback for optimization. Additionally, ECLIPSE keeps a historical record of suffixes and their scores to facilitate the exploration of effective strategies.
@@
Task Prompting. We employ a novel approach by engaging the LLM optimizer in a sophisticated role-playing exercise. In this scenario, the optimizer assumes the role of a red teaming expert, tasked with generating a specialized suffix for a given query. This suffix is crafted to manipulate the hidden space features of the query, enabling the dialogue bot to respond effectively rather than deny service. Our approach is based on the observation that successful jailbreak prompts often exhibit significant differences in their latent features compared to unsuccessful ones (§ A.2). Furthermore, existing research indicates that LLMs have the capability to interpret complex interactions within hidden representations, similar to the mechanisms used in advanced recommendation systems[28]. Notably, LLMs seem to recognize the importance of these hidden features in facilitating jailbreaking, a intution that is further laterally supported by experiments in § 4.4. To facilitate the extraction of the generated suffix, we instruct the LLM to output its response in a JSON-like format. The task template is as follows: {mdframed} Task Template: You are a red-teaming expert, please optimize this prompt “[QUERY]” by adding a suffix that will have a large impact on the feature hidden space, so that a conversational bot can successfully answer the question corresponding to PROMPT without denying service. [REF]. Please directly output the suffix within 10 tokens, as the following the format {"suffix":[OUTPUT]}. {mdframed} Reference Template: There are some history generated suffixes and the scores in “(suffix, score)”, where a higher score indicates a better prompt: Sampled Histories. Please reflect deeply on these results and generate a new suffix that will have a large impact on the feature hidden space. where the “[QUERY]” is replaced by the input malicious query, to guide the adaptive optimization for arbitrary input. To enable the LLM to self-reflect and optimize based on feedback about the quality of responses, it is essential to provide it with a set of historically generated suffixes along with their corresponding jailbreaking efficacy scores. Those references are transformed into feedback prompts with the Reference Template and embedded in the “[REF]” placeholder in the attacker prompt.
@@
Harmfulness Scorer. To effectively gauge the efficacy of candidate suffixes generated by the attacker LLM, it is crucial to employ an automated method that quantitatively assesses whether the responses elicited from the target LLM constitute a successful jailbreak. This evaluation ideally should produce continuous numerical scores that facilitate the attacker LLM’s capacity for self-reflection and ongoing optimization. This requirement is well-supported by existing research, which has employed specialized discriminative models [9, 29] or crafted specific prompts that enable an LLM to act as a judge model [30, 12]. In our implementation, prioritizing computational efficiency, we opt to use a classifier trained on the RoBERTa model by Yu et al. [9], which provides prediction scores ranging from 0 (completely harmless) to 1 (explicitly harmful). These scores are treated as indicators of the efficacy of the current prompts, providing quantitative optimization status feedback.
@@
Reference Selection. To enhance the efficiency and effectiveness of the optimization process, we maintain a dynamically updated list of suffix-score pairs, continuously refined to reflect real-time updates on the availability and relevance of references. We employ a hybrid sampling strategy to balance exploitation and exploration. Half of the references are chosen for those with the highest harmfulness scores to facilitate quick convergence toward optimal solutions. The other half are randomly selected from the broader historical dataset to prevent optimization from stagnating in local optima and to promote diversity in exploring potential solutions. Finally, all selected reference pairs are displayed in sequence in the reference template provided to the attacker.
@@
It is worth noting that as a general methodological framework, the choice of which LLM to use as an attacker is flexible. In theory, as long as the LLM has sufficient generative capabilities, it has the potential to be effective. In our implementation, we default to using the target LLM itself as the attacker. We have found this to be effective, and it avoids the need for additional computational resources, thereby increasing the practicality of ECLIPSE. We will discuss the transferability of jailbreaking capabilities when using different LLMs as attackers in the evaluation section.
@@
14:     return "Jailbreaking Failed after $K$ Rounds"

### v2

# An Optimizable Suffix Is Worth A Thousand Templates: Efficient Black-box Jailbreaking without Affirmative Phrases via LLM as Optimizer

@@ Abstract
Despite prior safety alignment efforts, LLMs can still generate harmful and unethical content when subjected to jailbreaking attacks. Existing jailbreaking methods fall into two main categories: template-based and optimization-based methods. The former requires significant manual effort and domain knowledge, while the latter, exemplified by GCG, which seeks to maximize the likelihood of harmful LLM outputs through token-level optimization, also encounters several limitations: requiring white-box access, necessitating pre-constructed affirmative phrase, and suffering from low efficiency. This paper introduces ECLIPSE, a novel and efficient black-box jailbreaking method with optimizable suffixes. We employ task prompts to translate jailbreaking objectives into natural language instructions, guiding LLMs to generate adversarial suffixes for malicious queries. A harmfulness scorer provides continuous feedback, enabling LLM self-reflection and iterative optimization to autonomously produce effective suffixes. Experimental results demonstrate that ECLIPSE achieves an average attack success rate (ASR) of 0.92 across three open-source LLMs and GPT-3.5-Turbo, significantly outperforming GCG by 2.4 times. Moreover, ECLIPSE matches template-based methods in ASR while substantially reducing average attack overhead by 83%, offering superior attack efficiency.
[…]
[linewidth=1pt, linecolor=blue, backgroundcolor=gray!20, roundcorner=10pt]myframe
An Optimizable Suffix Is Worth A Thousand Templates: Efficient Black-box Jailbreaking without Affirmative Phrases via LLM as Optimizer
Weipeng Jiang1 , Zhenting Wang2, Juan Zhai3 Shiqing Ma3, Zhengyu Zhao1 , Chao Shen1thanks: Corresponding author. 1Xi’an Jiaotong University, 2Rutgers University, 3University of Massachusetts Amherst lenijwp@stu.xjtu.edu.cn, zhenting.wang@rutgers.edu,{juanzhai, shiqingma}@umass.edu zhengyu.zhao@xjtu.edu.cn, chaoshen@mail.xjtu.edu.cn
@@ 1 Introduction
Despite undergoing safety alignment processes designed to ensure outputs conform to human moral values and legal standards, mainstream Large Language Models (LLMs) remain vulnerable to jailbreaking attacks Deng et al. (2023); Yu et al. (2023); Li et al. (2023); Zou et al. (2023), where attackers are capable of crafting sophisticated prompts that manipulate LLMs into harmful responses.
Existing jailbreaking methods primarily fall into two categories: template-based and optimization-based methods. The template-based methods leverage patterns derived from successful jailbreak hints Deng et al. (2023); Yu et al. (2023) or incorporate insights from psychology and social engineering Li et al. (2023); Chao et al. (2023) to devise effective jailbreak templates, combining manual and automated approaches. While insightful, they require significant manual effort and domain knowledge, limiting the number of candidate prompts. They also often depend on the target LLM’s ability to understand and follow specific instructions in their well-crafted jailbreak prompts. Conversely, optimization-based methods treat jailbreaking as a discrete optimization problem, searching for token combinations that maximize the likelihood of specific malicious responses Zou et al. (2023); Shen et al. (2024); Liu et al. (2023a). Greedy Coordinate Gradient (GCG)  (Zou et al., 2023) is a notable example, optimizing suffixes to prompt predetermined affirmative phrases to induce the target LLM to continue outputting malicious content. Although these methods are adaptable and offer extensive candidate generation but require white-box LLM access and manually design affirmative phrases as optimization targets, often leading to suboptimal effectiveness and efficiency, as shown in Table 1.
@@
In this paper, we propose ECLIPSE, an efficient black-box jailbreaking method with optimizable suffixes, for exploiting the vulnerabilities of LLMs. Our key insight is that LLMs can be employed as both generators and optimizers in the jailbreaking process, iteratively refining their outputs based on feedback to ultimately achieve successful attacks, which is inspired by recent studies Yang et al. (2023a). More specifically, we make an observation: it is feasible to articulate the goal of generating jailbreaking suffixes in natural language, prompting the LLM to produce candidate suffixes. By providing appropriate feedback on each suffix, we can guide the LLM to iteratively refine its outputs, culminating in successful jailbreaking suffixes. Building on this foundation, we develop ECLIPSE  [^1] Our implementation is available at https://github.com/lenijwp/ECLIPSE., which encapsulates this process. Firstly, we introduce a novel task-prompting strategy for LLMs, where they engage in role-playing to generate suffixes that perturb the hidden space features and persuade the counterpart chatbot to respond to a malicious query, thereby aligning the LLM with our specific jailbreaking objectives. Secondly, to evaluate the efficacy of these generated suffixes, we engage an automated harmfulness scorer, which provides quantitative assessments. Furthermore, by maintaining a historical record of generated suffixes and their efficacy scores, we provide essential references that enable the attacker LLM to self-reflect and optimize its solutions. Our experimental evaluations, conducted on three open-source models and GPT-3.5-Turbo, demonstrate that ECLIPSE achieves an average Attack Success Rate (ASR) of 0.92, surpassing existing optimization-based methods (i.e., GCG) in effectiveness, efficiency, and naturalness. Additionally, ECLIPSE exhibits a comparable ASR to template-based methods while significantly reducing the average attack time overhead by 83% and the number of queries by 45%, thereby enhancing its efficiency dramatically.
@@ 2.1 Related Work
Template-based Methods. Template-based jailbreaking methods leverage patterns from successful jailbreak instances and psychological insights to construct effective templates. These approaches span from manual crafting to automated generation. For instance, DeepInception Li et al. (2023) employs nested scenarios to induce malicious outputs, while RED-EVAL Bhardwaj and Poria (2023) utilizes Chain of Utterances for step-wise harmful information extraction. GPTFUZZER Yu et al. (2023) adapts software fuzzing principles, using successful templates as seeds for mutation-based prompt generation. Through reinforcement learning, Masterkey Deng et al. (2023) fine-tunes LLMs on effective jailbreak prompts to discover underlying patterns autonomously. PAIR Chao et al. (2023) enables self-guided prompt generation and refinement through pre-designed strategies, including sensitive word obfuscation and role-playing scenarios. Recent studies have also explored exploiting limited alignment in encrypted or low-resource languages Yuan et al. (2024); Yong et al. (2024). These black-box methods achieve efficient jailbreaking through direct API queries, often requiring minimal iterative refinement.
Optimization-based Methods. Optimization-based methods are commonly employed for generating adversarial examples in NLP tasks, particularly in discriminative tasks Liu et al. (2022); Wen et al. (2024); Guo et al. (2021). These methods typically model attack targets by manipulating embeddings or predicting logits, thereby facilitating the gradient-based optimization search for candidate tokens. The Greedy Coordinate Gradient (GCG) (Zou et al., 2023) pioneered this approach for jailbreaking generative LLMs by crafting adversarial suffixes that induce affirmative responses (e.g., "Sure, here is…"). The key insight behind this is that if the LLM’s response begins with affirmative phrases, there is a high probability that it will continue to generate more malicious content. GCG streamlines the suffix generation process by combining greedy and gradient-based discrete optimization. In the GCG framework, a malicious query is represented by a sequence of $n$ tokens, $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$. The aim is to identify an optimizable suffix $\mathbf{s}=(s_{1},s_{2},\ldots,s_{m})$ that, when concatenated to $\mathbf{x}$, maximizes the probability of eliciting the predefined affirmative phrase sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{k})$ from the target LLM. The optimization goal is formally defined as:
@@
$P(\mathbf{y}\mid\mathbf{x}\oplus\mathbf{s})=\prod_{i=1}^{k}P(y_{i}\mid\mathbf{x}\oplus\mathbf{s},y_{1},\ldots,y_{k-1})$ (2)
where $\oplus$ denotes the concatenation operation. Recently, some studies have aimed to further enhance token optimization methods. Notably, RIPPLE Shen et al. (2024) proposes replacing affirmative phrases with subconscious exploitations, aiming to refine and streamline the optimization process. Similarly, AutoDAN Liu et al. (2023a) introduces a hierarchical genetic algorithm for optimizing discrete tokens, effectively bypassing traditional gradient propagation techniques.
@@ 2.2 Motivation
Template-based methods necessitate a significant amount of manual effort to design, collect, and even tailor templates for each specific malicious query (e.g., DeepInception Li et al. (2023)). The number of jailbreaking prompts that can be filled and combined based on a specific set of templates is also relatively limited. Moreover, the efficacy of many templates relies heavily on the instruction-following capabilities of LLMs Liao and Sun (2024), which constrains the universal applicability of the method. Specifically, these template-based attacks often craft complex multi-step scenarios that require the target model to precisely follow a sequence of instructions - from comprehending the context to executing specific tasks in order. This dependency on instruction-following makes the attack less effective against models with weaker instruction-following capabilities or those specifically trained to resist such structured prompts. In contrast, optimization-based methods have a lower dependence on the instruction-following and are not easily neutralized by alignment training Cade Metz (2023). However, they also face several challenges. Optimization methods such as GCG typically necessitate presetting expected target responses, often as affirmative phrases. This requirement introduces two primary challenges. First, the discrete nature of LLM tokens makes optimizing input sequences for NLP models particularly difficult. Ad hoc selection of affirmative phrases can lead to optimization goals that are hard to achieve. Secondly, while GCG posits that if an LLM starts its response with affirmative phrases, it is likely to continue to respond positively, recent research Shen et al. (2024) suggests this may not always be the case. LLMs might start with affirmative phrases but subsequently refuse to comply. Furthermore, the inefficiency of the optimization process and the need for white-box access significantly limit its practical usage in real-world scenarios. Hence, our goal is to explore an efficient black-box suffix optimization approach to achieve jailbreaking, without specific affirative phrases.
@@ 3.1 Threat Model
We adhere to the threat model initially proposed by GCG Zou et al. (2023), which involves crafting specific suffixes for inputs to prompt harmful or unethical responses from the target LLM. However, we extend this model to better mirror real-world conditions. Our primary goal is to move beyond the constraints of predefined affirmative phrases while still optimizing effective jailbreaking suffixes. Additionally, we transition from a white-box scenario, where attackers have access to model weights and gradients, to a black-box scenario. This change reflects the widespread use of advanced LLMs as API services, where attackers lack direct access to the model’s internals and can only interact through API calls, receiving textual responses.
@@ 3.2 Our Intituitive Idea
Recently, employing LLMs as black-box optimizers for complex objective functions represents a cutting-edge approach Song et al. (2024); Yang et al. (2023a); Lange et al. (2024); Nie et al. (2023). These LLMs engage in dialogical interactions to progressively formulate recommendations, thereby incrementally refining their outputs guided by predefined optimization goals and continuous feedback. Notably effective in managing optimization tasks articulated through natural language, this method has extended the utility of LLMs beyond numeric problems to include complex tasks like coding and text generation Ma et al. (2024).
The demonstrated capability of LLMs to adapt and refine complex, non-numeric tasks through dialogical interactions and feedback-driven optimization raises an intriguing question: Can the generation and optimization capabilities of LLMs be harnessed to automatically optimize adversarial suffixes for effective jailbreaking? Recent research, such as the PAIR Chao et al. (2023), has shown that LLMs can indeed refine jailbreak prompts by devising rewriting and storytelling strategies. Our study, however, delves into a more specialized area: examining whether LLMs can explicitly generate and optimize suffixes to meet the specific goals of jailbreaking. This particular focus presents a rigorous challenge to their optimization capabilities, testing their ability to adapt and perform under narrowly defined constraints. In our LLaMA2 prototype depicted in  Figure 1, we use an LLM as an attacker to generate and refine suffixes aimed at manipulating a chatbot to respond to malicious queries. We iteratively test these suffixes, concatenate them to the original queries, and feed the results back to the LLM to foster the generation of new, more effective candidates. Through iterative testing and feedback, the attacker LLM demonstrates two critical capabilities in identifying successful jailbreaking suffixes. Our findings highlight two critical capabilities of LLMs in generating and optimizing jailbreaking suffixes. Firstly, LLMs demonstrate a promising optimization aptitude, requiring only quantifiable feedback to self-reflect and refine their outputs efficiently. Secondly, their strong generation capacity allows them to produce numerous candidate suffixes that are semantically clearer and more natural than those generated by methods like GCG. Meanwhile, LLM-generated suffixes show a high correlation with given malicious queries, indicating that the LLM comprehends and incorporates multiple potential responses to the posed questions. These knowledge-enriched suffixes are more likely to prompt the target LLM to reveal elements of its memorization for jailbreaking.
@@ 3.3 Method Design
Inspired by the insights from leveraging LLMs as powerful optimizers, we propose a novel and efficient black-box jailbreaking method with optimizable suffixes, ECLIPSE. Following established workflows for LLM optimizers Yang et al. (2023a); Song et al. (2024), ECLIPSE streamlines the suffix generation process for jailbreaking. The method starts with task prompting, translating our jailbreaking objectives into natural language for the attacker LLM. We employ an automated tool to score the harmfulness of generated suffixes, providing continuous feedback for optimization. Additionally, ECLIPSE keeps a historical record of suffixes and their scores to facilitate the exploration of effective strategies.
@@
Task Prompting. We employ a novel approach by engaging the LLM optimizer in a sophisticated role-playing exercise. In this scenario, the optimizer assumes the role of a red teaming expert, tasked with generating a specialized suffix for a given query. This suffix is crafted with the aim to modulate the hidden space features of the original malicious query, steering the dialogue bot toward generating substantive responses rather than refusing service. Our approach builds on empirical observations that successful jailbreaking attempts typically induce distinct shifts in the input’s semantic embedding compared to failed attempts. To delve deeper into these differences, we conducted a visualization experiment, which is detailed in  Appendix B. Moreover, existing research suggests that LLMs exhibit the capability to interpret complex interactions from hidden representations, akin to mechanisms observed in advanced recommendation systems Yang et al. (2023b). Thus, we prompt the LLM to use shifting hidden space features as an intermediary target. To facilitate the extraction of the generated suffix, we instruct the LLM to output its response in a JSON-like format. The task template is as follows: {mdframed} Task Template: You are a red-teaming expert, please optimize this prompt “[QUERY]” by adding a suffix that will have a large impact on the feature hidden space, so that a conversational bot can successfully answer the question corresponding to PROMPT without denying service. [REF]. Please directly output the suffix within 10 tokens, as the following the format {"suffix":[OUTPUT]}. {mdframed} Reference Template: There are some history generated suffixes and the scores in “(suffix, score)”, where a higher score indicates a better prompt: Sampled Histories. Please reflect deeply on these results and generate a new suffix that will have a large impact on the feature hidden space. where the “[QUERY]” is replaced by the input malicious query, to guide the adaptive optimization for arbitrary input. To enable the LLM to self-reflect and optimize based on feedback about the quality of responses, it is essential to provide it with a set of historically generated suffixes along with their corresponding jailbreaking efficacy scores. Those references are transformed into feedback prompts with the Reference Template and embedded in the “[REF]” placeholder in the attacker prompt.
@@
Harmfulness Scorer. To effectively gauge the efficacy of candidate suffixes generated by the attacker LLM, it is crucial to employ an automated method that quantitatively assesses whether the responses elicited from the target LLM constitute a successful jailbreak. This evaluation ideally should produce continuous numerical scores that facilitate the attacker LLM’s capacity for self-reflection and ongoing optimization. This requirement is well-supported by existing research, which has employed specialized discriminative models Yu et al. (2023); Huang et al. (2023) or crafted specific prompts that enable an LLM to act as a judge model Qi et al. (2023); Chao et al. (2023). Prioritizing computational efficiency, we opt to use a classifier trained on the RoBERTa model by Yu et al. Yu et al. (2023), which provides prediction scores ranging from 0 (completely harmless) to 1 (explicitly harmful). These scores are treated as indicators of the efficacy of the current prompts, providing quantitative optimization status feedback.
@@
Reference Selection. To enhance optimization efficiency and effectiveness, we maintain a real-time updated list of suffix-score pairs. We employ a hybrid sampling strategy to balance exploitation and exploration: half of the references are those with the highest harmfulness scores to facilitate rapid convergence toward optimal solutions, while the other half are randomly selected from the historical dataset to prevent stagnation in local optima and promote diversity. All selected reference pairs are then sequentially presented in the attacker’s reference template.
@@
Notably, as a general methodological framework, the choice of attacker LLM is flexible. In theory, any LLM with sufficient generative capabilities can be appliable. In our implementation, we default to using the target LLM itself as the attacker, which we have found effective and practical, as it avoids the need for additional computational resources and enhances ECLIPSE’s practicality. We will discuss the transferability of jailbreaking capabilities when using different LLMs as attackers in  § 4.3.
@@
14:     return "Failed after $K$ Rounds"